# Learning with Local Search MCMC Layers

### v1

@@ 3.2 Mixing neighborhood systems
Central to local search algorithms in combinatorial optimization is the use of multiple neighborhood systems to more effectively explore the solution space [37, 10]. In this section, we propose an efficient way to incorporate such diversity of neighborhood systems into the combinatorial MCMC layer, while preserving the correct stationary distribution, by mixing the corresponding proposal distributions.
@@
We assume that the individual Metropolis correction ratios $\tilde{\alpha}_{s}(\bm{y},\bm{y}^{\prime})\coloneqq\frac{q_{s}(\bm{y^{\prime}},\bm{y})}{q_{s}(\bm{y},\bm{y}^{\prime})}$ are tractable. The proposed procedure is summarized in Algorithm 2.
@@
If each neighborhood graph $G_{{\mathcal{N}}_{s}}$ is undirected and without self-loops, and the aggregate neighborhood graph $G_{\bar{\mathcal{N}}}$ is connected, the iterations ${\bm{y}^{(k)}}$ produced by Algorithm 2 follow a Markov chain with unique stationary distribution ${\pi}_{\bm{\theta},t}$.
See Section C.3 for the proof. Importantly, only the connectedness of the aggregate neighborhood graph $G_{\bar{\mathcal{N}}}$ is required. This allows us to combine neighborhood systems that could not connect $\mathcal{Y}$ if used individually, i.e., an irreducible Markov chain can be obtained by mixing the proposal distributions of reducible ones. Such an example is given with the moves defined in Table 2.
@@ 4 Loss functions and theoretical analysis
In this section, we derive and study loss functions for learning models using the proposed layer.
@@ 4.1 Negative log-likelihood and associated Fenchel-Young loss
We now show that the proposed layer ${\widehat{\bm{y}}_{t}}({\bm{\theta}})$ can be viewed as the solution of a regularized optimization problem on $\mathcal{C}=\operatorname{\mathsf{conv}}(\mathcal{Y})$. Let $A_{t}({\bm{\theta}})\coloneqq t\cdot\log\sum_{\bm{y}\in\mathcal{Y}}\exp\left(\left[\langle\bm{\theta}\,,\,\bm{y}\rangle+\varphi(\bm{y})\right]/t\right)$ be the cumulant function [53] associated to the exponential family defined by ${\pi}_{\bm{\theta},t}$, scaled by the temperature $t$. We define the regularization function $\Omega_{t}$ and the corresponding Fenchel-Young loss [8] as:
@@
Since $\Omega_{t}=A_{t}^{*}$ is strictly convex on $\operatorname{\mathsf{relint}}(\mathcal{C})$ (see Section C.4 for a proof) and ${\widehat{\bm{y}}_{t}}({\bm{\theta}})=\nabla_{{\bm{\theta}}}A_{t}({\bm{\theta}})$, the proposed layer is the solution of the regularized optimization problem
${\widehat{\bm{y}}_{t}}({\bm{\theta}})=\operatorname*{\mathsf{argmax}}_{\bm{\mu}\in\cal{C}}\left\{\langle\bm{\theta},\bm{\mu}\rangle-\Omega_{t}(\bm{\mu})\right\},$
the Fenchel-Young loss $\ell_{t}$ is differentiable, satisfies $\ell_{t}({\bm{\theta}},{\bm{y}})=0\Leftrightarrow{\widehat{\bm{y}}_{t}}({\bm{\theta}})={\bm{y}}$, and has gradient $\nabla_{\bm{\theta}}\ell_{t}(\bm{\theta}\,;\bm{y})={\widehat{\bm{y}}_{t}}({\bm{\theta}})-\bm{y}$. It is therefore equivalent, up to a constant, to the negative log-likelihood loss, as we have $-\nabla_{{\bm{\theta}}}\log\pi_{{\bm{\theta}},t}({\bm{y}})=({\widehat{\bm{y}}_{t}}({\bm{\theta}})-{\bm{y}})/t$. Algorithms 1 and 2 can thus be used to perform maximum likelihood estimation, by returning a (biased) stochastic estimate of the gradient of $\ell_{t}$.
@@ 4.2 Empirical risk minimization
In the supervised learning setting, we are given observations $\left({\bm{x}}_{i},\,{\bm{y}}_{i}\right)_{i=1}^{N}\in\left(\mathbb{R}^{p}\times\mathcal{Y}\right)^{N}$, and want to fit a model $g_{W}\colon{\mathbb{R}}^{p}\to{\mathbb{R}}^{d}$ such that ${\widehat{\bm{y}}_{t}}(g_{W}(\bm{x}_{i}))\approx{\bm{y}}_{i}$. This is motivated by a generative model where, for some weights $W_{0}\in{\mathbb{R}}^{p}$, the data is generated with ${\bm{y}}_{i}\sim{\pi}_{g_{W_{0}}(\bm{x}_{i}),t}$. We aim at minimizing the empirical risk $L_{N}$, defined below along with its exact gradient $\nabla L_{N}$ :
$L_{N}(W)\!\coloneqq\!\frac{1}{N}\!\sum_{i=1}^{N}\ell_{t}\left(g_{W}(\bm{x}_{i});{\bm{y}}_{i}\right)\quad\text{and}\quad\nabla_{W}L_{N}(W)\!=\!\frac{1}{N}\!\sum_{i=1}^{N}J_{W}g_{W}(\bm{x}_{i})\left({\widehat{\bm{y}}_{t}}(g_{W}(\bm{x}_{i}))\!-\!{\bm{y}}_{i}\right).$
@@
In practice, we cannot compute the exact gradient above. Using Algorithm 1 to get a MCMC estimate of ${\widehat{\bm{y}}_{t}}(g_{W}(\bm{x}_{i}))$, we propose the following estimator:
$\nabla_{W}L_{N}(W)\approx J_{W}g_{W}(\bm{x}_{i})\left(\frac{1}{K}\sum_{k=1}^{K}\bm{y}_{i}^{(k)}\!\!-{\bm{y}}_{i}\right),$
where $\bm{y}_{i}^{(k)}$ is the $k$-th iterate of Algorithm 1 with maximization direction $\bm{\theta}_{i}=g_{W}(\bm{x}_{i})$ and temperature $t$. This estimator is doubly stochastic, since we sample both data points and iterations of Algorithm 1, and can be seamlessly used with batches. The term $J_{W}g_{W}(\bm{x}_{i})$ is computed via autodiff.
@@
Following the contrastive divergence literature [22], in the supervised setting, we initialize the Markov chains at the corresponding ground-truth structure, by setting ${\bm{y}}_{i}^{(0)}={\bm{y}}_{i}$. In the unsupervised setting, we use a persistent initialization method [50] instead.
@@ 4.3 Associated Fenchel-Young loss with a single MCMC iteration
In Section 4.1, we saw that there exists a Fenchel-Young loss (the negative log-likelihood loss) associated with the proposed layer ${\widehat{\bm{y}}_{t}}$. However, it assumes that the number of MCMC iterations goes to infinity ($K\to\infty$). In practice, since we always use a finite number of iterations $K$, this means that our doubly stochastic gradient estimator is biased with respect to that loss.
We now show that when a single MCMC iteration is used ($K=1$), there exists another target-dependent Fenchel-Young loss such that the stochastic gradient estimator is unbiased with respect to that loss. See Section C.7 for the construction of $\Omega_{{\bm{y}}}$ and the proof. {proposition}[Existence of a Fenchel-Young loss when $K=1$] Let $\bm{p}^{(1)}_{\bm{\theta},{\bm{y}}}$ denote the distribution of the first iterate of the Markov chain defined by the Markov transition kernel given in Eq. 3, with proposal distribution $q$ and initialized at ground-truth ${\bm{y}}\in\mathcal{Y}$. There exists a target-dependent regularization function $\Omega_{{\bm{y}}}$ with the following properties: $\Omega_{{\bm{y}}}$ is $t/\mathbb{E}_{q({\bm{y}},\,\cdot\,)}||Y-{\bm{y}}||_{2}^{2}$-strongly convex; it is such that
$\mathbb{E}_{\bm{p}^{(1)}_{\bm{\theta},{\bm{y}}}}\left[Y\right]=\operatorname*{\mathsf{argmax}}_{\bm{\mu}\in\operatorname{\mathsf{conv}}\left(\mathcal{N}({\bm{y}})\cup\{{\bm{y}}\}\right)}\left\{\langle\bm{\theta},\bm{\mu}\rangle-\Omega_{{\bm{y}}}(\bm{\mu})\right\}$
and the Fenchel-Young loss $\ell_{\Omega_{{\bm{y}}}}$ generated by $\Omega_{{\bm{y}}}$ is $\mathbb{E}_{q({\bm{y}},\,\cdot\,)}||Y-{\bm{y}}||_{2}^{2}/t$-smooth in its first argument, and such that $\nabla_{\bm{\theta}}\ell_{\Omega_{{\bm{y}}}}(\bm{\theta}\,;{\bm{y}})=\mathbb{E}_{\bm{p}^{(1)}_{\bm{\theta},{\bm{y}}}}\left[Y\right]-{\bm{y}}$. A similar result in the unsupervised setting with data-based initialization is given in Section B.1. Interestingly, theses results contrast with prior work on the expected CD-$1$ update. Indeed, when applied with Gibbs sampling to train restricted Boltzmann machines, the latter was shown in Sutskever and Tieleman [49] not to be the gradient of any function – let alone a convex one.
4.4 Convergence analysis in the unsupervised setting
In the unsupervised setting, we are only given observations $\left({\bm{y}}_{i}\right)_{i=1}^{N}\in\mathcal{Y}^{N}$ and want to fit a model ${\pi}_{\bm{\theta},t}$, motivated by an underlying generative model such that ${\bm{y}}_{i}\sim{\pi}_{\bm{\theta}_{0},t}$ for an unknown true parameter $\bm{\theta}_{0}$. We assume here that $\mathcal{C}=\operatorname{\mathsf{conv}}(\mathcal{Y})$ is of full dimension in $\mathbb{R}^{d}$. If not, such a model is specified only up to vectors orthogonal to the direction of the smallest affine subspace spanned by $\mathcal{C}$. Indeed, if $\bm{\mu}\in\mathbb{R}^{d}$ is such that $\forall\,\bm{y}\in\mathcal{Y},\,\langle\bm{\mu},\,\bm{y}\rangle=C\in\mathbb{R}$, then we have ${\pi}_{\bm{\theta},t}={\pi}_{\bm{\theta}+\bm{\mu},t}$.
We have the corresponding empirical $L_{N}$ and population $L_{{\bm{\theta}}_{0}}$ Fenchel-Young losses:
which are minimized for $\bm{\theta}$ such that ${\widehat{\bm{y}}_{t}}(\bm{\theta})=\bar{Y}_{N}\coloneqq\frac{1}{N}\sum_{i=1}^{N}\bm{y}_{i}$, and for $\bm{\theta}$ such that ${\widehat{\bm{y}}_{t}}(\bm{\theta})={\widehat{\bm{y}}_{t}}(\bm{\theta}_{0})$, respectively. Let $\bm{\theta}^{\star}_{N}$ as the minimizer of the empirical loss $L_{N}$. For it to be defined, we assume $N$ is large enough to have $\bar{Y}_{N}\in\operatorname{\mathsf{relint}}(\mathcal{C})$ (which is always possible as ${\pi}_{\bm{\theta}_{0},t}$ has dense support on $\mathcal{Y}$). A slight variation on Proposition 4.1 in Berthet et al. [6], proved in Section C.5, gives the following asymptotic normality as $N\to\infty$. {proposition}[Convergence of the empirical loss minmizer to the true parameter]
We now consider the sample size as fixed to $N$ samples, and define $\hat{\bm{\theta}}_{n}$ as the $n$-th iterate of the following stochastic gradient algorithm:
where $\bm{y}^{(n+1,k)}$ is the $k$-th iterate of Algorithm 1 with temperature $t$, maximization direction $\hat{\bm{\theta}}_{n}$, and initialized at $\bm{y}^{(n+1,1)}=\bm{y}^{(n,K_{n})}$. This initialization corresponds to the persistent contrastive divergences (PCD) algorithm [50], and is further discussed in Section B.2.
{proposition}
[Convergence of the stochastic gradient estimate] Suppose the following assumptions on the step sizes $\left(\gamma_{n}\right)_{n\geq 1}$, sample sizes $\left(K_{n}\right)_{n\geq 1}$, and proposal distribution $q$ hold:
@@ 4.4 Convergence analysis in the unsupervised setting
  • $K_{n+1}>\left\lfloor 1+a^{\prime}\exp\left(\frac{8R_{\mathcal{C}}}{t}\cdot||\hat{\bm{\theta}}_{n}||\right)\right\rfloor$ with $a^{\prime}>0$ and $R_{\mathcal{C}}=\max_{\bm{y}\in\mathcal{Y}}||\bm{y}||$,
  • $\frac{1}{\sqrt{K_{n}}}-\frac{1}{\sqrt{K_{n-1}}}\leq a^{\prime\prime}n^{-c}$, with $a^{\prime\prime}>0$ and $c>1-\frac{b}{2}$,
  • $q(\bm{y},\bm{y}^{\prime})=\left\{\begin{array}[]{ll}\frac{1}{2d^{*}}&\text{if $\bm{y}^{\prime}\in\mathcal{N}(\bm{y})$,}\\ 1-\frac{d(\bm{y})}{2d^{*}}&\text{if $\bm{y}^{\prime}=\bm{y}$,}\\ 0&\text{else},\end{array}\right.$
    where $d(\bm{y})\coloneqq|\mathcal{N}(\bm{y})|$ is the degree of $\bm{y}$ in $G_{\mathcal{N}}$, and $d^{*}\coloneqq\max_{\bm{y}\in\mathcal{Y}}d(\bm{y})$.
@@
5 Experiments on dynamic vehicle routing
Overview of the challenge.
We evaluate the proposed approach on a large-scale, ML-enriched combinatorial optimization problem: the EURO Meets NeurIPS 2022 Vehicle Routing Competition [27]. In this dynamic vehicle routing problem with time windows (DVRPTW), requests arrive continuously throughout a planning horizon, which is partitioned into a series of delivery waves $\mathcal{W}=\{\left[\tau_{0},\tau_{1}\right],\left[\tau_{1},\tau_{2}\right],\dots,\left[\tau_{|\mathcal{W}|-1},\tau_{|\mathcal{W}|}\right]\}$.
$\min_{\Psi}\mathbb{E}\left[c_{\mathcal{W}}(f)\right],\quad\text{with}\quad c_{\mathcal{W}}(f)\coloneqq\sum_{\omega\in\mathcal{W}}c(f(\bm{x}^{\omega})),$
[…]
Reduction to supervised learning.
We follow the method of [4], which was the winning approach for the challenge. Central to this approach is the concept of prize-collecting dynamic vehicle routing problem with time windows (PC-VRPTW). In this setting, each request $i\in\mathcal{R}^{\omega}$ is assigned an artificial prize $\theta_{i}^{\omega}\in\mathbb{R}$, that reflects the benefit of serving it. The prize of the depot $D$ is set to $\theta^{\omega}_{D}=0$. The objective is then to identify a set of routes that maximizes the total prize collected while minimizing the associated travel costs. The model $g_{W}$ predicts the prize vector $\bm{\theta}^{\omega}=g_{W}(\bm{x}^{\omega})$. Denoting $\varphi(\bm{y})\coloneqq-\langle\bm{c},\bm{y}\rangle$, the corresponding optimization problem can be written as
The overall pipeline is summarized in Fig. 1. Following [4], we approximately solve the problem in Eq. 7 using the prize-collecting HGS heuristic (PC-HGS), a variant of hybrid genetic search (HGS) [51]. We denote this approximate solver $\widetilde{\bm{y}}\approx\widehat{\bm{y}}$, so that their proposed policy decomposes as $f_{W}\coloneqq\widetilde{{\bm{y}}}\circ g_{W}$. The ground-truth routes are created by using an anticipative strategy, i.e., by solving multiple instances where all future information is revealed from the start, and the requests’ arrival times information is translated into time windows (thus removing the dynamic aspect of the problem). This anticipative policy, which we note $f^{\star}$ (which cannot be attained as it needs unavailable information) is thus the target policy imitated by the model – see Section B.4 for more details.
Perturbation-based baseline.
In [4], a perturbation-based method [6] was used. This method is based on injecting noise in the PC-HGS solver $\widetilde{\bm{y}}$. Similarly to our approach, the parameters $W$ can then be learned using a Fenchel-Young loss, since the loss is minimized when the perturbed solver correctly predicts the ground truth. However, since $\widetilde{\bm{y}}$ is not an exact solver, all theoretical learning guarantees associated with this method (e.g., correctness of the gradients) no longer hold.
Proposed approach.
Our proposed approach instead uses the Fenchel-Young loss associated with the proposed layer, which is minimized when the proposed layer correctly predicts the ground-truth. At inference time, however, we use $f_{W}\coloneqq\widetilde{\bm{y}}\circ g_{W}$. We use a mixture of proposals, as defined in Algorithm 2. To design each proposal $q_{s}$, we build randomized versions of moves specifically designed for the prize-collecting dynamic vehicle routing problem with time windows. More precisely, we base our proposals on moves used in the local search part of the PC-HGS algorithm, which are summarized in Table 2. The details of turning these moves into proposal distributions with tractable individual correction ratios are given in Section B.3.
@@ 5 Experiments on dynamic vehicle routing
Performance metric.

### v2

@@ 3.2 Mixing neighborhood systems
Central to local search algorithms in combinatorial optimization is the use of multiple neighborhood systems to more effectively explore the solution space (Mladenović and Hansen, 1997; Blum and Roli, 2003). In this section, we propose a tractable way to incorporate such diversity of neighborhood systems into the combinatorial MCMC layer, while preserving the correct stationary distribution, by mixing the corresponding proposal distributions.
Definitions.
@@
Mixing proposals.
A naive way to combine these neighborhood systems would be to use Algorithm˜1 by defining an aggregated proposal $q({\bm{y}},\,\cdot\,)$, with support $\bar{{\mathcal{N}}}({\bm{y}})$ or $\bar{{\mathcal{N}}}({\bm{y}})\cup\{{\bm{y}}\}$, as:
$q({\bm{y}},{\bm{y}}^{\prime})\coloneqq\frac{1}{|Q({\bm{y}})|}\sum_{s\in Q({\bm{y}})}q_{s}({\bm{y}},{\bm{y}}^{\prime}),\quad\text{giving:}\quad\alpha({\bm{y}},{\bm{y}}^{\prime})=\frac{|Q({\bm{y}})|}{|Q({\bm{y}}^{\prime})|}\cdot\frac{\sum_{s\in Q({\bm{y}}^{\prime})}q_{s}({\bm{y}}^{\prime},{\bm{y}})}{\sum_{s\in Q({\bm{y}})}q_{s}({\bm{y}},{\bm{y}}^{\prime})}.$
However, this leads to intractable updates due to the computation of the correction ratio. Indeed, computing $\alpha({\bm{y}},{\bm{y}}^{\prime})$ is prohibitively expensive as it involves summing the forward proposal probabilities for all move types in $Q({\bm{y}})$ and the reverse probabilities for all move types in $Q({\bm{y}}^{\prime})$. The difficulty is that multiple, distinct proposal types can generate the same solution $\bm{y}$$^{\prime}$ from $\bm{y}$}. For example, in our vehicle routing application in Section˜5.1, relocating a pair of clients (see the relocate pair move from Table˜2) before the first one in a route of three gives the same solution $\bm{y}$$^{\prime}$ as relocating the first client (with relocate) at the last position. Identifying and calculating all these potential forward and reverse pathways for every step is a significant computational hurdle.
In contrast, the update we propose in Algorithm˜2 only requires computing the single individual ratio $\alpha_{s}({\bm{y}},{\bm{y}}^{\prime})\coloneqq\frac{|Q({\bm{y}})|}{|Q({\bm{y}}^{\prime})|}\cdot\frac{q_{s}(\bm{y}^{\prime},\bm{y})}{q_{s}(\bm{y},\bm{y}^{\prime})}$ for the unique move type $s$ that was actually sampled.
@@
If each neighborhood graph $G_{{\mathcal{N}}_{s}}$ is undirected and without self-loops, and the aggregate neighborhood graph $G_{\bar{\mathcal{N}}}$ is connected, the iterations $({\bm{y}^{(k)}})_{k\in\mathbb{N}}$ produced by Algorithm˜2 follow a Markov chain with unique stationary distribution ${\pi}_{\bm{\theta},t}$.
See Section˜D.3 for the proof. Importantly, only the aggregate neighborhood graph $G_{\bar{\mathcal{N}}}$ is required to be connected. This enables combining neighborhood systems ${\mathcal{N}}_{s}$ that could not connect $\mathcal{Y}$ if used individually, and an irreducible Markov chain can be obtained by mixing the proposal distributions of reducible ones. As a concrete example, the moves used as proposals in our dynamic vehicle routing experiment (Section˜5.1) are defined in Table˜2.
@@ 4 Loss functions and theoretical analysis
Building upon the differentiable MCMC layer developed in Section 3, this section constructs the corresponding learning framework. We derive principled Fenchel-Young loss functions for our layer, present practical stochastic gradient algorithms for both conditional and unconditional learning, and provide theoretical convergence guarantees for these algorithms.
@@ 4.1 Negative log-likelihood and associated Fenchel-Young loss
We now show that the proposed layer ${\widehat{\bm{y}}_{t}}({\bm{\theta}})$ can be viewed as the solution of a regularized optimization problem on $\mathcal{C}=\operatorname{\mathsf{conv}}(\mathcal{Y})$. Let $A_{t}({\bm{\theta}})\coloneqq t\cdot\log\sum_{\bm{y}\in\mathcal{Y}}\exp\left(\left[\langle\bm{\theta}\,,\,\bm{y}\rangle+\varphi(\bm{y})\right]/t\right)$ be the cumulant function (Wainwright and Jordan, 2008) associated to ${\pi}_{\bm{\theta},t}$, scaled by $t$. We define the regularization function $\Omega_{t}$ and the corresponding Fenchel-Young loss (Blondel et al., 2020) as:
@@
Since $\Omega_{t}=A_{t}^{*}$ is strictly convex on $\operatorname{\mathsf{relint}}(\mathcal{C})$ (see Section˜D.4 for a proof) and ${\widehat{\bm{y}}_{t}}({\bm{\theta}})=\nabla_{{\bm{\theta}}}A_{t}({\bm{\theta}})$, the proposed layer is the solution of the regularized optimization problem
${\widehat{\bm{y}}_{t}}({\bm{\theta}})=\operatorname*{\mathsf{argmax}}_{\bm{\mu}\in\cal{C}}\left\{\langle\bm{\theta},\bm{\mu}\rangle-\Omega_{t}(\bm{\mu})\right\},$ (6)
the Fenchel-Young loss $\ell_{t}$ is differentiable, satisfies $\ell_{t}({\bm{\theta}},{\bm{y}})=0\Leftrightarrow{\widehat{\bm{y}}_{t}}({\bm{\theta}})={\bm{y}}$, and has gradient $\nabla_{\bm{\theta}}\ell_{t}(\bm{\theta}\,;\bm{y})={\widehat{\bm{y}}_{t}}({\bm{\theta}})-\bm{y}$ (Blondel et al., 2020). It is therefore equivalent, up to a constant, to the negative log-likelihood loss, as we have $\nabla_{{\bm{\theta}}}\log\pi_{{\bm{\theta}},t}({\bm{y}})=({\widehat{\bm{y}}_{t}}({\bm{\theta}})-{\bm{y}})/t$. Algorithms 1 and 2 can thus be used to perform MLE, by returning a (biased) stochastic estimate of the gradient of $\ell_{t}$.
@@ 4.2 Empirical risk minimization
In the conditional learning setting, we are given observations $\left({\bm{x}}_{i},\,{\bm{y}}_{i}\right)_{i=1}^{N}\in\left(\mathbb{R}^{p}\times\mathcal{Y}\right)^{N}$, and want to fit a model $g_{W}\colon{\mathbb{R}}^{p}\to{\mathbb{R}}^{d}$ such that ${\widehat{\bm{y}}_{t}}(g_{W}(\bm{x}_{i}))\approx{\bm{y}}_{i}$. This is motivated by a generative model where, for some weights $W_{0}\in{\mathbb{R}}^{p}$, the data is generated with ${\bm{y}}_{i}\sim{\pi}_{g_{W_{0}}(\bm{x}_{i}),t}$. We aim at minimizing the empirical risk $L_{N}$, defined below along with its exact gradient:
$L_{N}(W)\!\coloneqq\!\frac{1}{N}\!\sum_{i=1}^{N}\ell_{t}\left(g_{W}(\bm{x}_{i})\,;{\bm{y}}_{i}\right),\;\;\text{with}\quad\nabla_{W}L_{N}(W)\!=\!\frac{1}{N}\!\sum_{i=1}^{N}J_{W}g_{W}(\bm{x}_{i})\left({\widehat{\bm{y}}_{t}}(g_{W}(\bm{x}_{i}))\!-\!{\bm{y}}_{i}\right).$
@@
In practice, we cannot compute the exact gradient above. Using Algorithm˜1 or 2 to get a MCMC estimate of ${\widehat{\bm{y}}_{t}}(g_{W}(\bm{x}_{i}))$, we propose the following estimator:
$\nabla_{W}L_{N}(W)\approx J_{W}g_{W}(\bm{x}_{i})\biggl(\frac{1}{K}\sum_{k=1}^{K}\bm{y}_{i}^{(k)}\!\!-{\bm{y}}_{i}\biggr),$
where $\bm{y}_{i}^{(k)}$ is the $k$-th iterate of the algorithm with maximization direction $\bm{\theta}_{i}=g_{W}(\bm{x}_{i})$ and temperature $t$. This estimator is doubly stochastic, since we sample both data points and Markov iterations, and can be seamlessly used with batches. The term $J_{W}g_{W}(\bm{x}_{i})$ is computed via autodiff.
@@
Following the contrastive divergence literature (Hinton, 2000), in the conditional setting, we initialize the Markov chains at the corresponding ground-truth, by setting ${\bm{y}}_{i}^{(0)}={\bm{y}}_{i}$. In the unconditional setting, we use a persistent initialization (Tieleman, 2008) instead.
@@ 4.3 Associated Fenchel-Young loss with a single MCMC iteration
To obtain an unbiased gradient estimator for the Fenchel-Young loss $\ell_{t}$ associated with $\widehat{\bm{y}}_{t}$}, the MCMC sampler must be run until it reaches its stationary distribution $\pi_{{\bm{\theta}},t}$. This requirement makes any practical estimator with a finite number of steps $K$ inherently biased.
@@
$\mathbb{E}_{\bm{p}^{(1)}_{\bm{\theta},{\bm{y}}}}\left[Y\right]=\operatorname*{\mathsf{argmax}}_{\bm{\mu}\in\operatorname{\mathsf{conv}}\left(\mathcal{N}({\bm{y}})\cup\{{\bm{y}}\}\right)}\left\{\langle\bm{\theta},\bm{\mu}\rangle-\Omega_{{\bm{y}}}(\bm{\mu})\right\},$
@@
A similar result in the unconditional setting with data-based initialization is given in Section˜B.1. In contrast, Sutskever and Tieleman (2010) showed that the expected CD-1 update with Gibbs sampling for restricted Boltzmann machines is not the gradient of any function, let alone a convex one.
@@
4.4 Convergence analysis in the unconditional setting
In the unconditional setting, we are given observations $\left({\bm{y}}_{i}\right)_{i=1}^{N}\!\!\in\!\!\mathcal{Y}^{N}\!$ and want to fit a model ${\pi}_{\bm{\theta},t}$, motivated by an underlying generative model such that ${\bm{y}}_{i}\sim{\pi}_{\bm{\theta}_{0},t}$ for some true parameter $\bm{\theta}_{0}$. We assume here that $\mathcal{C}=\operatorname{\mathsf{conv}}(\mathcal{Y})$ is of full dimension in $\mathbb{R}^{d}$ (if not, the model is specified only up to vectors $\bm{\mu}$} orthogonal to the affine subspace spanned by $\mathcal{C}$, as $\pi_{{\bm{\theta}}+{\bm{\mu}},t}\!=\!\pi_{{\bm{\theta}},t}$). We have the corresponding empirical $L_{N}$ and population $L_{{\bm{\theta}}_{0}}$ Fenchel-Young losses:
which are minimized for $\bm{\theta}$ such that ${\widehat{\bm{y}}_{t}}(\bm{\theta})=\bar{Y}_{N}\coloneqq\frac{1}{N}\sum_{i=1}^{N}\bm{y}_{i}$, and for $\bm{\theta}$ such that ${\widehat{\bm{y}}_{t}}(\bm{\theta})={\widehat{\bm{y}}_{t}}(\bm{\theta}_{0})$, respectively. Let $\bm{\theta}^{\star}_{N}$ as the minimizer of the empirical loss $L_{N}$. For it to be defined, we assume that $\bar{Y}_{N}\in\operatorname{\mathsf{int}}(\mathcal{C})$ (which is always the case for $N$ large enough, as ${\pi}_{\bm{\theta}_{0},t}$ has dense support on $\mathcal{Y}$). A slight variation on Proposition 4.1 in Berthet et al. (2020) gives the following asymptotic normality: {proposition}[Convergence of the empirical loss minimizer to the true parameter]
The proof is given in Section˜D.5. We now consider the sample size as fixed to $N$ samples, and define $\hat{\bm{\theta}}_{n}$ as the $n$-th iterate of the following stochastic gradient algorithm:
where $\bm{y}^{(n+1,k)}$ is the $k$-th iterate of Algorithm˜1 with temperature $t$, maximization direction $\hat{\bm{\theta}}_{n}$, and initialized at $\bm{y}^{(n+1,1)}=\bm{y}^{(n,K_{n})}$. This initialization corresponds to the persistent contrastive divergences (PCD) algorithm (Tieleman, 2008), and is further discussed in Section˜B.3. {proposition}[Convergence of the stochastic gradient estimate] Suppose the following hold for the step sizes $(\gamma_{n})_{n\geq 1}$, sample sizes $(K_{n})_{n\geq 1}$, and proposal distribution $q$:
$\gamma_{n}=an^{-b}$, with $b\in\left(\tfrac{1}{2},1\right]$ and $a>0$.
$K_{n+1}>\lfloor 1+a^{\prime}\exp(\tfrac{8R_{\mathcal{C}}}{t}\|\hat{\bm{\theta}}_{n}\|)\rfloor$, with $a^{\prime}>0$ and $R_{\mathcal{C}}=\max_{\bm{y}\in\mathcal{Y}}\|\bm{y}\|$.
$\tfrac{1}{\sqrt{K_{n}}}-\tfrac{1}{\sqrt{K_{n-1}}}\leq a^{\prime\prime}n^{-c}$, with $a^{\prime\prime}>0$ and $c>1-\tfrac{b}{2}$.
$q(\bm{y},\bm{y}^{\prime})=\begin{cases}\tfrac{1}{2d^{*}},&\bm{y}^{\prime}\in\mathcal{N}(\bm{y}),\\ 1-\tfrac{d(\bm{y})}{2d^{*}},&\bm{y}^{\prime}=\bm{y},\\ 0,&\text{else},\end{cases}$ where $d(\bm{y})\coloneqq|\mathcal{N}(\bm{y})|$ and $d^{*}\coloneqq\max_{\bm{y}\in\mathcal{Y}}d(\bm{y})$.
Then the iterates $\hat{\bm{\theta}}_{n}$ defined by Eq.˜7 converge almost surely: $\hat{\bm{\theta}}_{n}\xrightarrow{\textit{a.s.}}\bm{\theta}^{\star}_{N}$. See Section˜D.6 for the proof. The assumptions on $q$ are used for obtaining a closed-form convergence rate bound for the Markov chain, using graph-based geometric bounds (Ingrassia, 1994).
5 Numerical experiments
5.1 Dynamic vehicle routing
We empirically validate our approach on the dynamic vehicle routing problem with time windows (DVRPTW) from the EURO Meets NeurIPS 2022 Vehicle Routing Competition (Kool et al., 2023). A detailed introduction to the challenge with precise notations is given in Appendix˜C.
Reduction to supervised learning.
In this DVRPTW, requests arrive in delivery waves $\omega$, at the start of which a dispatching and vehicle routing problem for the current set of requests $\cal{R}^{\omega}$ must be solved, to get a feasible solution ${\bm{y}}^{\omega}\in\mathcal{Y}(\cal{R}^{\omega})$. Following Baty et al. (2023), we frame each dispatching and routing problem as a prize-collecting (PC-)VRP, where a model $g_{W}$ predicts a "prize" vector ${\bm{\theta}}^{\omega}$ for serving each request. This PC-VRP fits the general formulation of Eq.˜2:
$\widehat{{\bm{y}}}({\bm{\theta}}^{\omega})=\operatorname*{\mathsf{argmax}}_{{\bm{y}}\in\mathcal{Y}(\cal{R}^{\omega})}\langle{\bm{\theta}}^{\omega},{\bm{y}}\rangle+\varphi({\bm{y}}),$ (8)
where $\varphi({\bm{y}})\coloneqq-\langle\bm{c},{\bm{y}}\rangle$ is the negative routing cost. The overall pipeline is shown in Fig.˜1. The model is trained to imitate an anticipative oracle $f^{\star}$, i.e., we use its output as ground-truth for supervised learning. We compute $f^{\star}$ by solving a static VRPTW where all future information in the instance is revealed from the start, turning dispatching waves into time windows.
[…]
Approach and baseline.
The baseline Baty et al. (2023), winner of the competition, relies on a perturbation-based method (Berthet et al., 2020) with the state-of-the-art PC-HGS heuristic $\tilde{{\bm{y}}}$ (Vidal, 2022) as a combinatorial optimization layer. Since $\tilde{{\bm{y}}}$ is an inexact solver, the theoretical guarantees granted by the framework of Berthet et al. (2020) no longer hold. Our approach instead uses a local search MCMC layer to train $g_{W}$. We use a mixture of proposals (Algorithm˜2) defined precisely in Section˜B.4, derived from the local search moves used by the PC-HGS solver itself (which are summarized in Table˜2). At inference time, we follow the baseline, and use $f_{W}\coloneqq\tilde{{\bm{y}}}\circ g_{W}$.